OpenStudy (anonymous):

∫_0^4▒(4e^5+ 3x)dx evaluate ..

OpenStudy (anonymous):

\[∫_0^4(4e^5+ 3x)dx \]

OpenStudy (anonymous):

if you can explain also that would be good

OpenStudy (anonymous):

\[4e^5x + \frac{3}{2}x^2 |_0^4\]

OpenStudy (anonymous):

how did you get 3/2timesx^2 with the 4/0

OpenStudy (anonymous):

x is x^2/2

OpenStudy (anonymous):

so do you understand it now

OpenStudy (anonymous):

i have no clue what i am doing here thats why i am asking

OpenStudy (anonymous):

kk i will review on my own will post few more

OpenStudy (anonymous):

how do you evaluate this further

OpenStudy (anonymous):

4e5x+32x2|40\[4e^{5}x+32x^{2}|\left(\begin{matrix}4 \\ 0\end{matrix}\right)\]

OpenStudy (anonymous):

what is with these binomials?

OpenStudy (anonymous):

@av8188 your job is to find a function whose derivative is between the integral sign and the dx once you do that you evaluate at the upper limit, then evaluate at the lower limit, and then subtract the lower from the upper

OpenStudy (anonymous):

also i am going to bet that you meant \[\int_0^4 (4e^{5x}+3x)dx\] eys?

OpenStudy (anonymous):

e^5x no it is not there

OpenStudy (anonymous):

so you need to find the "anti derivative of \[4e^{5x}\] which is \[\frac{4}{5}e^{5x}\] and the anti derivative of \[3x\] which is \[\frac{3}{2}x^2\] in other words \[\int (4e^{3x}+3)dx=\frac{4}{5}e^{5x}+\frac{3}{2}x^2\]

OpenStudy (anonymous):

then plug in 4, plug in 0, and subtract

OpenStudy (anonymous):

hey how did you type 4/0

OpenStudy (anonymous):

you mean \[\int_0^4\]

OpenStudy (anonymous):

use \int_0^4

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

but we need to be clear about the first term. i am betting it is \[4e^{5x}\] and not \[4e^5\] which one is right?

OpenStudy (anonymous):

look at ishans post the end of it has |4/0

OpenStudy (anonymous):

no sorry

OpenStudy (anonymous):

|_0^4

OpenStudy (anonymous):

this thread on the top

OpenStudy (anonymous):

that will do it

OpenStudy (anonymous):

ah k

OpenStudy (anonymous):

so how to solve this now sorry to bug u

OpenStudy (anonymous):

underscore gives subscript

OpenStudy (anonymous):

happy to do it but let me repeat. what is the first term?

OpenStudy (anonymous):

to be honest man i dont know what i am doing here

OpenStudy (anonymous):

yes and i can walk you through no problem

OpenStudy (anonymous):

but just so we don't get some wrong answer i still need to know what he first term is is it \[4e^5\] or is it \[4e^{5x}\]?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4e^{5}

OpenStudy (anonymous):

\[4e^{5}\]

OpenStudy (anonymous):

ok then it is easy. because \[4e^5\] is a constant

OpenStudy (anonymous):

want to go slow or just get answer?

OpenStudy (anonymous):

answer is fine on this one i have 2 more so on those you can explain

OpenStudy (anonymous):

if answer is all you need it is here http://www.wolframalpha.com/input/?i=integral+0+to+4+%284e^5%2B3x%29+dx

OpenStudy (anonymous):

but this is a very easy one, steps would be simple

OpenStudy (anonymous):

please show steps lol

OpenStudy (anonymous):

ok here we go

OpenStudy (anonymous):

is there a site that shows the steps? save you time...

OpenStudy (anonymous):

first of all it is clear that \[4e^5\] is just a number, a constant, not a function. and you want \[\int_0^4 4e^5 dx\] meaning the area under the curve (line) \[y=4e^5\] from x = 0 to x = 4. this is just a rectangle, whose base if 4 (from 0 to 4 is 4) and whose height is \[4e^5\] so the area of a rectangle is just base times height, we get \[\int_0^4 4e^5dx=4\times 4e^5=16e^5\]

OpenStudy (anonymous):

now we want \[\int_0^43xdx\] and here you have to find a function whose derivative is 3x that function is \[\frac{3}{2}x^2\] and that should be clear once you see it

OpenStudy (anonymous):

k

OpenStudy (anonymous):

then you can write \[\frac{3}{2}x^2|_0^4\] if yo like but this just means plug in 4, plug in 0 and subtract. if you plug in 4 you get \[\frac{3}{2}4^2=\frac{3}{2}\times 16=3\times 8=24\] and if you plug in 0 you get 0

OpenStudy (anonymous):

so in total we have \[\int_0^4 4e^5+3xdx=16e^5+24\] finished

OpenStudy (anonymous):

just like worlfram said in the link i sent you

OpenStudy (anonymous):

thank you

OpenStudy (zarkon):

you probably already know, but I'll show anyway for \[\frac{3}{2}x^2|_0^4\] it looks nicer when you do this \[\left.\frac{3}{2}x^2\right|_0^4\] \left.\frac{3}{2}x^2\right|_0^4

OpenStudy (anonymous):

actually i didn't know. thanks. do i need the left? or just the right?

OpenStudy (zarkon):

you need both

OpenStudy (anonymous):

\[\frac{3}{2}x^2\right|_0^4\]

OpenStudy (anonymous):

yeah i guess so

OpenStudy (anonymous):

\[\left.x^3 \right|_0^4\]

OpenStudy (anonymous):

oh and i need the dot too. didn't see that one

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

\[\left.\frac{5}{4}x^3\right|_0^5\]

OpenStudy (zarkon):

it just makes the vertical line the same size as you expression

OpenStudy (anonymous):

aah it expands to fit the expression i see

OpenStudy (anonymous):

thanks!

OpenStudy (zarkon):

just like the parentheses

OpenStudy (zarkon):

\[\left(\frac{3}{4}\right)\]

OpenStudy (zarkon):

\left(\frac{3}{4}\right)

OpenStudy (anonymous):

damn mine didn't work

OpenStudy (anonymous):

\[\left(\frac{1}{\frac{1}{1+x}}\right)^2\]

OpenStudy (zarkon):

\[f(x)=\left\{\begin{matrix}x^2, & x>0 \\ 2-x, & x\leq0\end{matrix}\right.\] f(x)=\left\{\begin{matrix}x^2, & x>0 \\ 2-x, & x\leq0\end{matrix}\right.

OpenStudy (anonymous):

now it did. no dot on this one

OpenStudy (anonymous):

wow i was doing this with array!

OpenStudy (zarkon):

only use the dot if you don't want to show anything

OpenStudy (anonymous):

ooh i see

OpenStudy (anonymous):

no screwing upo

OpenStudy (zarkon):

sorry that didn't copy correctly f(x)=\left\{\begin{matrix}x^2, & x>0 \\ 2-x, & x\leq0\end{matrix}\right. hmmm...this is not displaying correctly...I am using the matrix command \begin{matrix}

OpenStudy (anonymous):

can you give me the whole thing you wrote for piecewise

OpenStudy (zarkon):

I'll have to give it in pieces

OpenStudy (anonymous):

if you don't mind just copy it in chat

OpenStudy (anonymous):

that way i can see it whole because latex does not work in chat

OpenStudy (anonymous):

f(x)=\left\{\begin{matrix}x^2, & x>0 \\ 2-x, & x\leq0\end{matrix}\right.

OpenStudy (anonymous):

\[f(x)=\left\{\begin{matrix}x^2, & x>0 \\ 2-x, & x\leq0\end{matrix}\right.\]

OpenStudy (zarkon):

nice :)

OpenStudy (anonymous):

damn damn damn

OpenStudy (anonymous):

i went back to chat and it was in latex there too!

OpenStudy (anonymous):

latex works in chat when you have question opened matt told me that

OpenStudy (anonymous):

do me a favor and send it again

OpenStudy (anonymous):

go to group home and then check the latex

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

pasted in to text editor

OpenStudy (anonymous):

i was using this f(x) = \left\{ \begin{array}{lr} 1 & : x \in \mathbb{Q}\\ 0 & : x \notin \mathbb{Q} \end{array} \right.

OpenStudy (zarkon):

ic

OpenStudy (zarkon):

you know you can see the source code for latex by right clicking on the latex and pressing on 'show source'?

OpenStudy (anonymous):

\[f(x)=\left\{\begin{matrix}x^2, & x<0 \\ 2-x, & 0<x<1\\e^x,&x\geq 1\end{matrix}\right.\]

OpenStudy (zarkon):

nice :)

OpenStudy (anonymous):

man i learn something new every day

OpenStudy (anonymous):

and all this time i have been rewriting when i could just copy and past from the code dope slap for me

OpenStudy (anonymous):

\[f(x)=\left\{\begin{matrix}x^2, & x<0 \\ 2-x, & 0<x<1\\e^x,&x\geq 1\end{matrix}\right.\]